OpenStudy (anonymous):

Express radical 48 in simplest form

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \sqrt{48}}\)\(\Huge\color{blue}{ \sf =\sqrt{3} \times \sqrt{16}}\)\(\Huge\color{blue}{ \sf =\sqrt{3} \times 4 }\)\(\Huge\color{blue}{ \sf =4\sqrt{3} }\)